the curve y=ax2+bx+c passes the points (1,8), (0,5) and (3,20)
find the values of a,b,c

Answer by Alan3354(69443)

the curve y=ax2+bx+c passes the points (1,8), (0,5) and (3,20)
find the values of a,b,c
------------
Sub for x and y for each of the 3 points, then solve the system of 3 equations for a, b and c.
-----------
Or, use determinants.
6y - 6x^2- 12x = 30
6y = 6x^2 + 12x + 30
y = x^2 + 2x + 5